OpenStudy (anonymous):

Decide whether the function is an exponential growth or exponential decay function and find the constant percentage rate of growth or decay. f(x) = 9.8 ⋅ 1.03x   Exponential growth function; 103% Exponential growth function; 0.03% Exponential growth function; 3% Exponential decay function; 103%

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

I think its C not 100% sure

OpenStudy (michele_laino):

an exponential growth function is like below: \[\huge y = A \cdot {B^x}\] where \(B>1\)

OpenStudy (michele_laino):

and the percentage rate\(r\) of growth, is defined as follows: \[\Large r = \frac{{y\left( {x + 1} \right) - y\left( x \right)}}{{y\left( x \right)}} = \frac{{A \cdot {B^{x + 1}} - A \cdot {B^x}}}{{A \cdot {B^x}}}=...?\] please complete, furthermore, keep in mind that \(B=1.03\)

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

it is \(f(x) = 9.8 ⋅ 1.03^x\) right?

OpenStudy (anonymous):

correct

OpenStudy (misty1212):

oh yeah, it is C it is always C

OpenStudy (anonymous):

is she correct? @Michele_Laino